# Floor Plan-Guided Visual Navigation Incorporating Depth and Directional Cues

### v2

@@ Abstract
Guiding an agent to a specific target in indoor environments based solely on RGB inputs and a floor plan is a promising yet challenging problem. Although existing methods have made significant progress, two challenges remain unresolved. First, the modality gap between egocentric RGB observations and the floor plan hinders the integration of visual and spatial information for both local obstacle avoidance and global planning. Second, accurate localization is critical for navigation performance, but remains challenging at deployment in unseen environments due to the lack of explicit geometric alignment between RGB inputs and floor plans. We propose a novel diffusion-based policy, denoted as GlocDiff, which integrates global path planning from the floor plan with local depth-aware features derived from RGB observations. The floor plan offers explicit global guidance, while the depth features provide implicit geometric cues, collectively enabling precise prediction of optimal navigation directions and robust obstacle avoidance. Moreover, GlocDiff introduces noise perturbation during training to enhance robustness against pose estimation errors, and we find that combining this with a relatively stable Visual Odometry (VO) module during inference results in significantly improved navigation performance. Extensive experiments on the FloNa benchmark demonstrate GlocDiff’s efficiency and effectiveness in achieving superior navigation performance, and the success of real-world deployments also highlights its potential for widespread practical applications.
Introduction
@@
To address this limitation, recent work (li2021cognitive; ewe2024spatial; li2024flona; setalaphruk2003robot; ericson2024beyond) has increasingly explored integrating prior floor plans with visual inputs. Floor plans constitute a valuable and widely available source of global spatial knowledge, providing high-level semantic and geometric layouts that remain invariant to dynamic changes such as human activity, temporary obstacles, or furniture rearrangement. This makes them well-suited as priors for supporting localization and navigation in unfamiliar environments. Despite their promise, existing methods still face two key challenges in fully utilizing the structural cues provided by floor plans to guide navigation effectively.
First, due to the inherent modality inconsistency between RGB observations and floor plans, effectively leveraging complementary clues from these two sources to provide local obstacle avoidance and global path planning guidance remains challenging. While prior approaches bypass this challenge by establishing extra topological or semantic maps, combining them with perceptions from LiDAR (ewe2024spatial; li2021cognitive) and inertial measurement units (goswami2024floor), they also introduce increased system costs. FloDiff (li2024flona) directly fuses latent features from floor plans and visual observations using transformer-based encoders, and performs end-to-end navigation within a diffusion policy framework. However, this straightforward fusion approach suffers from low data utilization efficiency and fails to exploit the critical geometric clues inherent in the inputs effectively.
Second, prior work (li2021cognitive; li2024flona) has shown that the accuracy of agent localization on the floor plan significantly affects navigation performance. Meanwhile, the spatial geometry misalignment between visual observations and floor plans makes achieving precise localization particularly challenging. Existing studies (chen2024f3loc; boniardi2019robot; howard2021lalaloc) tackle this challenge by matching geometric features from visual observations with boundary features of floor plans. However, these methods often struggle in environments with complex furniture arrangements or repetitive layouts, leading to inaccurate localization.
To address the above challenges, we propose GlocDiff, a diffusion-based policy for floor plan-guided visual navigation.The core idea of GlocDiff lies in leveraging depth and directional cues from RGB observations and the floor plan, respectively, to guide action generation. As shown in Fig. 1, GlocDiff compute the shortest path from the agent’s current location to the goal on the floor plan, which—despite potential intersections with obstacles—serves as a corse yet informative directional prior. Additionally, GlocDiff leverages a pre-trained model (ke2024repurposing) to produce depth features from RGB observations, enabling a better understanding of local scene geometry and identifying traversable regions, which enables safe navigation in complex environments. The path and depth guidance explicitly exploit geometric clues essential for next-step planning, thus bridging the modality gap inherent in the inputs and improving data efficiency.
Given the reliance of navigation on high-precision localization, we further improve GlocDiff in two ways. (i) Unlike previous studies, we employ a standard visual feature-based VO algorithm (freda2025pySLAM) as the localization module, ensuring more stable and reliable localization throughout the navigation process. (ii) We introduce noise into the current pose and path planning results during training to enhance the model’s robustness to localization inaccuracies.
@@
Extensive experiments on the FloNa (li2024flona) benchmark demonstrate that our approach achieves superior performance even with fewer training pairs, highlighting both its effectiveness and data efficiency. We also conduct comprehensive ablation studies to assess the impact of conditioning on the global shortest path and local depth features, as well as GlocDiff’s robustness to varying localization accuracy. Finally, we validate the practical applicability of GlocDiff by deploying it on an Automated Guided Vehicle (AGV) without any fine-tuning, showcasing its potential for real-world deployment in embodied systems.
Our contributions are summarized as follows: (1) We propose conditioning the navigation policy on global path planning and local depth features to provide more effective guidance, thus enabling efficient and collision-free navigation. (2) We deploy an online Visual Odometry (VO) system to ensure localization stability and enhance the policy’s robustness by injecting noise into the current pose and shortest path during training. (3) Extensive quantitative and qualitative experiments demonstrate that our method achieves state-of-the-art performance on the FloNa benchmark, and the ablation studies highlight the effectiveness of the proposed designs.
Related Work
Mapping-based Navigation
Mapping-based navigation, the prevailing framework, requires building an environmental map first to support subsequent planning and localization. Typically, VSLAM techniques (mur2015orb; mur2017orb) rely on visual observation inputs to construct sparse metric maps of the surrounding environment. In recent years, more researchers have integrated semantic information into map representations. Previous studies (cha2020obj; ram2022pon; guo2024object; yu2024trajectory) extract semantic categories from observations, associate the categories with each point in the point cloud, and finally generate a semantic map via differentiable projection operations and a denoising network. Other approaches integrate point-level features (huang2023visual), image features (pan2025planning; chen2022think; wang2023dreamwalker), and panoramic images (chaplot2020neural) into the built topological graph. VLFM (yokoyama2024vlfm) iteratively constructs an occupancy map and a value map to guide the agent toward the out-of-view target object. However, the mapping-based method has to explore unseen areas, which tends to be time-consuming for navigation. Moreover, the dynamic nature of objects in indoor environments necessitates real-time updates to the constructed map, making its reuse difficult and significantly reducing efficiency.
@@
Figure 2: Overview of GlocDiff. The depth encoder uses RGB observations to produce depth latent features, which are fed into a multi-head attention module, yielding the depth context. Taking the depth context, the current pose from the localization module, and the floor plan feature from the floor plan encoder as inputs, the MLP outputs the depth cue. The directional cue is obtained by computing the shortest path to the goal using the A algorithm. Conditioned on depth and directional cues, the diffusion-based policy generates future actions.
Mapless Navigation
@@
Without relying on mapping, mapless visual navigation methods directly focus on translating observations into actions. wijmans2020ddppo solve the point goal navigation with a near-perfect success rate under idealized settings through large-scale reinforcement learning. Learning-based visual odometry (partsey2022ismapping; zhao2021surprising) is used to address realistic point goal navigation. Recent work also leverages LLMs to guide agents in navigation (Zheng_2024_CVPR; chen2025affordances; zhou2024navgpt). Leveraging the diffusion policy (chi2023diffusion), NoMaD (sridhar2024nomad) produces collision-free actions and flexibly handles goal-conditioned and goal-agnostic navigation via the goal masking technique. Although mapless methods eliminate the need for mapping, they often make myopic decisions due to the lack of global navigable information.
Navigation with Floor Plan
@@
Compared to the two methods mentioned above, floor plan-based navigation eliminates the need for active exploration and instead utilizes the inherent structure and spatial information to perform reliable navigation actions. However, floor plans typically only depict permanent structures such as walls and lack information about obstacles like furniture or other objects. This absence poses challenges for collision-free planning. Setalaphruk et al. (setalaphruk2003robot) extract a Voronoi diagram from the sketch map and subsequently implement localization and navigation based on multiple hypotheses. li2021cognitive propose a cognitive navigation solution using floor plans. However, its dependency on pre-collected visual features hinders its practical applicability. To eliminate the dependency, Ewe et al. (ewe2024spatial) develop a graph-based navigation framework, enabling navigation without relying on precise scale information. goswami2024floor leverage stereo-inertial sensors to construct a 2D semantic point cloud, which, in conjunction with the floor plan, enables robot localization and subsequent navigation. These methods all utilize sensors other than monocular cameras. The closest related work to ours is FloNa (li2024flona), which marks the first attempt to navigate unseen environments using only a monocular camera and a floor plan. FloNa utilizes a transformer-based vision encoder to extract observation context from the floor plan and visual inputs. A diffusion policy (chi2023diffusion), conditioned on the context, is then employed to generate actions. Although FloNa demonstrates impressive performance, it fails to effectively integrate complementary geometric clues from observations and floor plans. GlocDiff tackles this issue by independently processing global and local information to derive two distinct types of guidance.
Diffusion Policy
@@
Diffusion-based policies have been successfully applied in various robotic domains, including dexterous manipulation (simeonov2023rpdiff; chi2023diffusion), locomotion (huang2024diffuseloco), and path planning (sridhar2024nomad; li2024flona). By framing robot behavior generation as a conditional denoising diffusion process, these policies not only offer high training stability but also ensure temporal consistency by generating coherent action sequences as a whole. Building upon these advances, our work focuses on utilizing diffusion policies conditioned on geometric clues derived from both visual observations and floor plans to achieve vision-based floor plan navigation.
Methodology
@@
Specifically, we employ a depth encoder (ke2024repurposing) to extract depth features from RGB images. This encoder utilizes a latent diffusion model to model the conditional distribution $P(\mathbf{d}|\mathbf{o})$, where $\mathbf{d}$ and $\mathbf{o}$ denote the low-dimensional depth feature and RGB observation, respectively. In our model, the depth encoder takes $5$ observations as input, including the current observation and the past $4$ observations, denoted as $\mathcal{O}=\{o_{t-4},o_{t-3},\ldots,o_{t}\}$. Then, the encoder outputs corresponding depth features denoted as $\mathcal{D}=\{d_{t-4},d_{t-3},\ldots,d_{t}\}$. These features are then passed through the multi-head attention module (vaswani2017attention), and the resulting fused features are averaged to derive the depth context. Subsequently, following (li2024flona), we employ a floor plan encoder consisting of an EfficientNet-B0 (tan2019efficientnet) network and a multilayer perceptron (MLP) to encode the floor plan image, obtaining a floor plan feature. Finally, the depth context, the floor plan feature, and the agent’s pose on the floor plan from the localization module are then concatenated and fed into an MLP to produce the local depth cue, $C_{l}$.
Localization Module
@@
During training, we use the ground truth (GT) poses from FloNa (li2024flona) as the localization results. For testing and deployment, we utilize a visual odometry or a floor plan-based visual localization module to estimate the current pose.
Directional Cue Module
@@
resulting in $\mathcal{T}_{a,s}=\{q_{a,0},q_{a,1},\cdots,q_{a,i},\cdots,q_{a,t}\}$. Here, $R=\bigl(\begin{smallmatrix}\cos{r_{0}}&-\sin{r_{0}}\\ \sin{r_{0}}&\cos{r_{0}}\end{smallmatrix}\bigr)$, $r_{0}$ denotes the orientation of the first point in the shortest path. In the agent coordinate system, the agent’s current position serves as the coordinate origin, the current orientation aligns with the positive X-axis, while the leftward direction aligns with the positive Y-axis. Notably, we assume the agent does not experience height variation during navigation, and thus, we only consider a 2D coordinate system (without the Z-axis). Finally, we empirically take the first $32$ points of $\mathcal{T}_{a,s}$ (about $1.2$ m) as the path guidance $C_{g}$.
Diffusion Policy
Diffusion policy (chi2023diffusion) leverages the Denoising Diffusion Probabilistic Models (DDPMs) (ho2020denoising) to formulate visuomotor robot policies. In our work, using the depth and directional cues, $C_{l}$ and $C_{g}$, the diffusion policy learns to approximate the conditional distribution $P(A_{t}|C_{l},C_{g})$ through a forward and reverse process, where $A_{t}$ is a $32\times 2$ vector representing the generated sequence of future $32$ actions at time step $t$. Each action corresponds to an approximate movement of $4$ cm, and the agent executes the first 20 steps (i.e., around $0.8$ cm) of the predicted sequence during deployment.
@@
In the reverse process, diffusion policy learns the conditional distribution $P(A_{t}^{k-1}|A_{t}^{k},C_{l},C_{g})$. Specifically, the diffusion policy employs a noise prediction network of UNet (ronneberger2015unet) architecture to predict the added noise, $\epsilon^{\prime}$, and utilizes the noise scheduler to carry out the denoising operation. The process can be described as:
@@
where $\varepsilon$ denotes the noise prediction network with parameters $\theta$ and $\mathcal{S}$ is the DDPM noise scheduler.
As shown in Fig. 3, our noise prediction network includes three downsample modules, one bottleneck module, two upsample modules, and a 1D convolutional layer. Each module consists of several 1D convolutional blocks, activation functions, and residual connections. The network takes noisy actions of size $32\times 2$, along with depth and path guidance, $C_{l}$ and $C_{g}$, as input and outputs a predicted noise of size $32\times 2$. $C_{l}$ and the time step serve as inputs to each module. Meanwhile, $C_{g}$ and time step are first processed through two identical residual blocks, producing two features: $f_{1}$ and $f_{2}$. $f_{1}$ is added to the output of the first residual block in the final upsample module, while $f_{2}$ is added to the output of the first residual block. Notably, we apply positional encoding to the time step before feeding it into the network. During training, we use the following loss function:
@@
Training
We use the data provided by FloNa (li2024flona), collected from $67$ Gibson simulation environments. Each scene contains $150$, $180$, or $200$ episodes, depending on the scene size, resulting in a total of $20,214$ episodes. We sample training data from this dataset, with each sample consisting of five consecutive frames, the current and goal positions, and a floor plan image.

### v3

@@ Abstract
Current visual navigation strategies mainly follow an exploration-first and then goal-directed navigation paradigm. This exploratory phase inevitably compromises the overall efficiency of navigation. Recent studies propose leveraging floor plans alongside RGB inputs to guide agents, aiming for rapid navigation without prior exploration or mapping. Key issues persist despite early successes. The modal gap and content misalignment between floor plans and RGB images necessitate an efficient approach to extract the most salient and complementary features from both for reliable navigation. Here, we propose GlocDiff, a novel framework that employs a diffusion-based policy to continuously predict future waypoints. This policy is conditioned on two complementary information streams: (1) local depth cues derived from the current RGB observation, and (2) global directional guidance extracted from the floor plan. The former handles immediate navigation safety by capturing surrounding geometry, while the latter ensures goal-directed efficiency by offering definitive directional cues. Extensive evaluations on the FloNa benchmark demonstrate that GlocDiff achieves superior efficiency and effectiveness. Furthermore, its successful deployment in real-world scenarios underscores its strong potential for broad practical application.
I Introduction
@@
To address this limitation, recent work [15, 7, 14, 25, 6] has increasingly explored integrating prior floor plans with visual inputs. Floor plans constitute a valuable and widely available source of global spatial knowledge, providing high-level semantic and geometric layouts that remain invariant to dynamic changes such as temporary obstacles or furniture rearrangement. This inherent stability makes them well-suited as priors for supporting navigation in unseen environments.
However, effectively integrating these two data sources for both local obstacle avoidance and global path planning remains challenging. The main bottleneck is the inherent mismatch between them: (1) a difference in the way spatial information is represented in egocentric observations versus abstract top-down architectural plans, and (2) a content gap, where objects (e.g., furniture) present in observations are missing from the static floor plan. While prior approaches bypass this challenge by establishing extra topological or semantic maps, combining them with perceptions from LiDAR [7, 15] and inertial measurement units [8], they also introduce increased system costs. Li et al. [14] employ a unified image encoder to extract latent features from both floor plans and visual observations, followed by transformer-based fusion for action prediction. This method achieves only a superficial and ineffective fusion. By processing the floor plan as a mere image, it squanders the inherent, explicit representation of global connectivity that is the defining feature of a human-abstracted topological map.
We propose GlocDiff, a diffusion-based policy for floor plan-guided visual navigation. The core idea of GlocDiff lies in leveraging depth and directional cues from RGB observations and the floor plan, respectively, to guide action generation. As shown in Fig. 1, GlocDiff bridges the representational gap by extracting a sequence of 2D waypoints (the shortest path) from the floor plan’s topology. This path—despite potential obstacle intersections—provides a geometrically grounded directional prior, translating the abstract map into actionable sub-goals for egocentric control. Concurrently, GlocDiff complements the global path with fine-grained local geometry. It leverages a depth encoder [13] to produce depth features from RGB observations, which reveal traversable regions and obstacles absent from the floor plan. To synthesize smooth and temporally coherent action sequences from these complementary signals, we employ a conditional diffusion model. It is conditioned on both guidance sources and trained to generate actions by supervised learning on optimal trajectories. To enhance robustness against the inevitable localization errors known to impact navigation performance [15, 14], we incorporate Gaussian noise perturbations into the agent’s pose and path planning results during training. Consequently, GlocDiff can be effectively deployed with a standard VO module [16], achieving robust performance even when faced with its noisy pose estimates.
@@
Extensive experiments on the FloNa [14] benchmark demonstrate that our approach outperforms the baselines even with fewer training pairs, emphasizing its effectiveness and data efficiency. We also conduct comprehensive ablation studies to assess the impact of conditioning on the global shortest path and local depth features, as well as GlocDiff’s robustness to varying localization accuracy. Finally, we validate the practical applicability of GlocDiff by deploying it on an Automated Guided Vehicle (AGV) without any fine-tuning, showcasing its potential for real-world deployment.
Our contributions are summarized as follows:
  • We propose conditioning the navigation policy on global path planning and local depth features to provide more effective guidance, thus enabling efficient and collision-free navigation.
  • Extensive experiments validate the effectiveness and generalization of our model. Additionally, the noise injection strategy during training ensures robust adaptation to VO, leading to successful deployment on the real robot.
@@
Figure 2: Overview of GlocDiff. The pre-trained depth encoder uses RGB observations to produce depth latent codes, which are fed into a multi-head attention module, yielding the depth context. Taking depth context, current pose from the localization module, and floor plan feature from the floor plan encoder as inputs, the MLP outputs geometry guidance. Path planner computes the shortest path from the current position to the goal using the A algorithm, serving as the path guidance. Conditioned on geometry and path guidance, the diffusion-based policy generates future actions.
II Related Work
II-A Mapping-based Navigation
Mapping-based navigation, the prevailing framework, requires building an environmental map first to support subsequent planning and localization. Typically, VSLAM techniques [17, 18] rely on visual observation inputs to construct sparse metric maps of the surrounding environment. In recent years, more works have started to integrate semantic information into map representations. Researchers [1, 21, 9] extract semantic categories from observations, associate the categories with each point in the point cloud, and finally generate a semantic map via differentiable projection operations and a denoising network. Other approaches integrate point-level features [11], image features [35, 4, 30], and panoramic images [2] into the built topological graph. VLFM [34] iteratively constructs an occupancy map and a value map to guide the agent toward the out-of-view target object. However, the mapping-based method has to explore unseen areas, which tends to be time-consuming for navigation. Moreover, the dynamic nature of objects in indoor environments necessitates real-time updates to the constructed map, making its reuse difficult and significantly reducing efficiency.
II-B Mapless Navigation
@@
Without relying on mapping, mapless visual navigation methods directly focus on translating observations into actions. Wjimans et al. [32] solve the point goal navigation with a near-perfect success rate under idealized settings through large-scale reinforcement learning. Learning-based visual odometry [20, 36] is used to address realistic point goal navigation. Wenzel et al. [31] train visual-motor policies for navigation tasks in a simple 3D environment. Work [23] design a lightweight perception network and convert visual navigation to scene classification. Leveraging the diffusion policy [5], NoMaD [27] produces collision-free future actions and flexibly handle both goal-conditioned and goal-agnostic navigation via the goal masking technique. Although mapless methods eliminate the need of mapping, they often produce shortsighted actions due to the lack of global navigable information.
II-C Navigation with Floor Plan
@@
Unlike the two methods mentioned above, floorplan-based navigation eliminates the need to explore the environment and instead utilizes the inherent structure and spatial information to perform reliable navigation actions. Setalaphruk et al. [25] extract a Voronoi diagram from the sketch map and subsequently implement localization and navigation based on multiple hypotheses. Li et al. [15] propose a cognitive navigation solution using floor plans. However, its dependency on pre-collect visual features hinders its practical applicability. To eliminate the dependency, Ewe et al. [7] develop a graph-based navigation framework, enabling navigation without relying on precise scale information. Goswami et al. [8] leverage stereo-inertial sensors to construct 2D semantic point cloud, which, in conjunction with the floor plan, enables robot localization and subsequent navigation. These methods all utilize sensors other than monocular cameras. The closest related work to ours is FloNa [14], which marks the first attempt to navigate unseen environments using only a monocular camera and a floor plan. FloNa utilizes a transformer-based vision encoder to extract observation context from the floor plan and visual inputs. A diffusion policy [5], conditioned on the context, is then employed to generate actions. Although FloNa demonstrates impressive performance, it fails to effectively integrate complementary geometric clues from observations and floor plans. GlocDiff tackles this issue by independently processing global and local information to derive two distinct guidances.
II-D Diffusion Policy
@@
Diffusion-based policies have been successfully applied in various robotic domains, including dexterous manipulation [26, 5], locomotion [12], and path planning [27, 14]. By framing robot behavior generation as a conditional denoising diffusion process, these policies not only offer high training stability but also demonstrate strong performance in modeling complex distributions. Building upon these advances, our work focuses on utilizing diffusion policies conditioned on geometric clues derived from both visual observations and floor plans to achieve vision-based floor plan navigation.
III Methodology
@@ III-A Depth Cue Module
Specifically, we employ a depth encoder [13] to extract depth features from RGB images. This encoder utilizes a latent diffusion model to model the conditional distribution $P(\mathbf{d}|\mathbf{o})$, where $\mathbf{d}$ and $\mathbf{o}$ denote the low-dimensional depth feature and RGB observation, respectively. In our model, the depth encoder takes $5$ observations as input, including the current observation and the past $4$ observations, denoted as $\mathcal{O}=\{o_{t-4},o_{t-3},\ldots,o_{t}\}$. Then, the encoder outputs corresponding depth features denoted as $\mathcal{D}=\{d_{t-4},d_{t-3},\ldots,d_{t}\}$. These features are then passed through the multi-head attention module [29], and the resulting fused features are averaged to derive the depth context. Subsequently, following [14], we employ a floor plan encoder consisting of an EfficientNet-B0 [28] network and a multilayer perceptron (MLP) to encode the floor plan image, obtaining a floor plan feature. Finally, the depth context, the floor plan feature, and the agent’s pose on the floor plan from the localization module (see Sec. III-B) are then concatenated and fed into an MLP to produce the local depth cue, $C_{l}$.
III-B Localization Module
@@
During training, we use the ground truth (GT) poses as the localization results. For testing and deployment, we utilize a visual odometry or a floor plan-based visual localization module to estimate the current pose.
III-C Directional Cue Module
@@
resulting in $\mathcal{T}_{a,s}=\{q_{a,0},q_{a,1},\cdots,q_{a,i},\cdots,q_{a,t}\}$. Here, $R=\bigl(\begin{smallmatrix}\cos{r_{0}}&-\sin{r_{0}}\\ \sin{r_{0}}&\cos{r_{0}}\end{smallmatrix}\bigr)$, $r_{0}$ denotes the orientation of the first point in the shortest path. In the agent coordinate system, the agent’s current position serves as the coordinate origin, the current orientation aligns with the positive X-axis, while the leftward direction aligns with the positive Y-axis. Notably, we assume the agent does not experience height variation during navigation, and thus, we only consider a 2D coordinate system (without the Z-axis). Finally, we take the first $m$ points of $\mathcal{T}_{a,s}$ (about $1.2$ m) as the path guidance $C_{g}$.
III-D Diffusion Policy
Diffusion policy [5] leverages the Denoising Diffusion Probabilistic Models (DDPMs) [10] paradigm to formulate visuomotor robot policies. In our work, using the depth and directional cues, $C_{l}$ and $C_{g}$, the diffusion policy learns to approximate the conditional distribution $P(A_{t}|C_{l},C_{g})$ through a forward and reverse process, where $A_{t}$ is a $m\times 2$ vector representing the generated sequence of future $m$ actions at time step $t$. Each action corresponds to an approximate movement of $4$ cm.
@@ III-D Diffusion Policy
In the reverse process, diffusion policy learns the conditional distribution $P(A_{t}^{k-1}|A_{t}^{k},C_{l},C_{g})$. Specifically, the diffusion policy employs a noise prediction network of UNet [24] architecture to predict the added noise, $\epsilon^{\prime}$, and utilizes the noise scheduler to carry out the denoising operation. The process can be described as:
@@
where $\varepsilon$ denotes the noise prediction network with parameters $\theta$ and $\mathcal{S}$ is the DDPM noise scheduler. As shown in Fig. 3, our noise prediction network includes three downsample modules, one bottleneck module, two upsample modules, and a 1D convolutional layer. Each module consists of several 1D convolutional blocks, activation functions, and residual connections. The network takes noisy actions of size $m\times 2$, along with depth and path guidance, $C_{l}$ and $C_{g}$, as input and outputs a predicted noise of size $m\times 2$. $C_{l}$ and the time step serve as inputs to each module. Meanwhile, $C_{g}$ and the time step are first processed through two identical residual blocks, producing two features: $f_{1}$ and $f_{2}$. $f_{1}$ is added to the output of the first residual block in the final upsample module, while $f_{2}$ is added to the output of the first residual block. Notably, we apply positional encoding to the time step before feeding it into the network. During training, we use the following loss function:
@@
TABLE I: SR ($\%$) and SPL ($\%$) of Different Models under Various Conditions. The Bold Number Indicates the Best Result.
Method Loc-A* (GT) DD-PPO (GT) Map-A* (GT) Loc-FloDiff (F3) Loc-FloDiff (GT) GlocDiff (F3) GlocDiff (VO) GlocDiff (GT)
Gibson HM3D Gibson HM3D Gibson HM3D Gibson HM3D Gibson HM3D Gibson HM3D Gibson HM3D Gibson HM3D
SR(%) $\uparrow$ $\tau_{c}$ $10$ $16.60$ $18.60$ $15.00$ $13.20$ $55.60$ $46.80$ $4.80$ $2.40$ $39.00$ $34.40$ $53.40$ $47.60$ $68.60$ $63.40$ $79.60$ $77.60$
$30$ $18.20$ $18.80$ $22.80$ $16.40$ $55.60$ $46.80$ $11.60$ $5.20$ $53.40$ $45.80$ $69.20$ $55.80$ $80.80$ $71.80$ $92.80$ $90.40$
$50$ $18.50$ $19.40$ $31.60$ $23.80$ $55.60$ $46.80$ $17.80$ $9.20$ $59.20$ $48.40$ $77.80$ $62.60$ $86.60$ $79.00$ $94.20$ $90.80$
$\infty$ $18.60$ $19.40$ $48.20$ $36.00$ $55.60$ $46.80$ $26.80$ $15.80$ $66.00$ $53.20$ $86.40$ $68.40$ $89.80$ $83.60$ $97.60$ $91.80$
[…]
$30$ $18.10$ $17.80$ $12.60$ $9.60$ $34.23$ $33.79$ $6.40$ $3.36$ $44.10$ $35.28$ $49.35$ $41.11$ $59.07$ $52.43$ $70.38$ $69.88$
$50$ $18.50$ $18.35$ $16.40$ $13.80$ $34.23$ $33.79$ $8.07$ $5.18$ $46.60$ $37.20$ $52.61$ $47.55$ $61.31$ $55.04$ $70.98$ $70.31$
$\infty$ $18.50$ $18.35$ $23.00$ $18.20$ $34.23$ $33.79$ $9.58$ $8.44$ $48.20$ $40.41$ $54.77$ $49.75$ $62.57$ $56.49$ $72.36$ $71.69$
III-E Training
We use the data provided by FloNa [14], collected from $67$ Gibson simulation environments. Each scene contains $150$, $180$, or $200$ episodes, depending on the scene size, resulting in a total of $20,214$ episodes. We sample training data from this dataset, with each sample consisting of a historical observation of five consecutive RGB frames, the agent’s current position, a goal position, a floor plan image, and the corresponding sequence of future waypoints. Unlike FloNa [14], which densely samples goal positions from future waypoints along the trajectory, we define the goal as the final waypoint of the episode. We sample one pair every five frames to avoid overlapping observations across different pairs. Consequently, our method is trained on 1.2 million samples, significantly fewer than the 180 million samples used for training Loc-FloDiff.